### v1

# Existence and large radial solutions for an elliptic system under finite new Keller-Osserman integral conditions

Abstract
We investigate the existence and qualitative behavior of entire positive radial solutions to the semilinear elliptic system
$\Delta u=p(|x|)\,g(v),\qquad\Delta v=q(|x|)\,f(u),\qquad x\in\mathbb{R}^{n},\ n\geq 3,$
under finite Keller–Osserman-type integral conditions on the nonlinearities $f$ and $g$, and integrability constraints on the radial weights $p$ and $q$. The nonlinearities are assumed continuous on $0,\infty)$, differentiable on $(0,\infty)$, vanish at the origin, and are strictly positive elsewhere, with
$\int_{1}^{\infty}\frac{dt}{g(f(t))}<\infty,\qquad\int_{1}^{\infty}\frac{dt}{f(g(t))}<\infty.$
The weights satisfy $\int_{0}^{\infty}s\,p(s)\,ds<\infty$, $\int_{0}^{\infty}s\,q(s)\,ds<\infty$, and $\min(p,q)$ is not compactly supported. Within this framework, we establish: (i) the existence of infinitely many entire positive radial solutions for a nonempty set of central values; (ii) closedness of the set of admissible central values; and (iii) largeness (blow-up at infinity) of solutions corresponding to boundary points of this set. The approach is based on a novel subharmonic functional tailored to the reciprocal integral conditions, extending classical Keller–Osserman theory to a broad class of coupled systems with general nonlinearities and weight functions.
MSC 2020: 35J47, 35B40, 35B09, 35J60.
Keywords and phrases: Semilinear elliptic systems, Keller–Osserman condition, radial solutions, large solutions, blow-up at infinity.
1 Introduction
The study of positive entire solutions to nonlinear elliptic systems has a long and rich history, driven both by deep theoretical challenges in nonlinear analysis and by diverse applications in physics, geometry, and biology. A classical starting point is the scalar equation
$\Delta u=f(u),$
for which Keller [5] and Osserman [6] established the celebrated Keller–Osserman condition, providing sharp integral criteria for the existence of large solutions (see also [11] fro more details). These foundational results have inspired extensive research on systems of elliptic equations, where the interaction between components introduces new analytical difficulties.
For semilinear elliptic systems, Lair and Wood [8] proved the existence of entire large positive solutions under suitable growth and integrability conditions, while Peterson-Wood [10] and Li-Yang [12] extended the analysis to non-monotone systems, but with superlinear nonlinearities. More recently, Lair [7] investigated competitive-type systems, obtaining sharp existence results for large solutions.
In addition to the aforementioned works, the contributions of [1, 3, 4, 13] have played a significant role in advancing the theory of elliptic systems, particularly in elucidating the crucial interplay between nonlinearity and weight functions.
In this work, we consider the semilinear elliptic system
$\Delta u=p(r)\,g(v),\qquad\Delta v=q(r)\,f(u),\qquad r=|x|,\quad x\in\mathbb{R}^{n},\ n\geq 3,$ (1)
under the following assumptions:
  • (F1)
    $f,g:[0,\infty)\rightarrow[0,\infty)$ are continuous on $0,\infty)$, differentiable on $(0,\infty)$, non-decreasing, with $f(0)=g(0)=0$ and $f(s),g(s)>0$ for all $s>0$;
  • (F2)
    multiplicative subadditivity:
    $f(sr)\leq f(s)f(r),\quad g(lt)\leq g(l)g(t),\quad\forall s,r,l,t\geq 0;$
  • (F3)
    finite Keller–Osserman-type integrals:
    $L_{f}:=\int_{1}^{\infty}\frac{dt}{\sqrt{\int_{0}^{t}g(f(z))\,dz}}<\infty,\quad L_{g}:=\int_{1}^{\infty}\frac{dt}{\sqrt{\int_{0}^{t}f(g(z))\,dz}}<\infty;$ (2)
  • (PQ)
    $p,q:[0,\infty)\rightarrow[0,\infty)$ are continuous, not both identically zero, and satisfy
    $\mathcal{L}_{p}$ $=\frac{1}{n-2}\int_{0}^{\infty}s\,p(s)\,ds=\lim_{r\rightarrow\infty}\mathcal{P}(r)<\infty,$
    $\mathcal{L}_{q}$ $=\frac{1}{n-2}\int_{0}^{\infty}s\,q(s)\,ds=\lim_{r\rightarrow\infty}\mathcal{Q}(r)<\infty,$ (3)
    where
    $\mathcal{P}(r)=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt,\quad\mathcal{Q}(r)=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt,$
    and $\min(p,q)$ does not have compact support.
We seek positive radial $C^{2}$-solutions $(u(r),v(r))$ of (1) with prescribed central values
$u(0)=a>0,\quad v(0)=b>0,\quad u^{\prime}(0)=v^{\prime}(0)=0.$
A positive entire large solution is a pair $(u,v)$ of $C^{2}$ radial functions such that $u(r),v(r)>0$ for all $r\geq 0$ and
$\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=\infty.$
In [7], Lair proved existence of such solutions for power-type nonlinearities $f(u)=u^{\theta}$, $g(v)=v^{\sigma}$ under certain integrability conditions on $p$ and $q$. Our goal is to extend this to general nonlinearities satisfying (2).
Remark 1 (Relation to Lair’s result).
When $g(v)=v^{\sigma}$ and $f(u)=u^{\theta}$ with $\sigma,\theta>0$ and $\sigma\cdot\theta>1$, the Keller–Osserman integrals in (2) converge, recovering the power-growth case in [7].
Our main objectives are:
  • to prove the existence of infinitely many positive entire radial solutions for a nonempty set of central values;
  • to show that the set of admissible central values is closed;
  • to establish that solutions corresponding to boundary points of this set are large, i.e., both components blow up at infinity.
The novelty of our approach lies in the construction of a new subharmonic functional adapted to the finite reciprocal integral conditions on $f$ and $g$, which allows us to complete, extend and unify the frameworks developed in [3, 4, 1, 5, 6, 8, 7, 10] to a broader class of nonlinearities and weight functions.
We now state our three main results.
Theorem 1.1 (Existence).
For every $(a,b)$ in a suitable nonempty set of admissible central values $\mathcal{T}\subset(0,\infty)^{2}$, there exists an entire positive radial solution $(u,v)\in C^{2}([0,\infty))^{2}$ of (1) with $u(0)=a$, $v(0)=b$. Moreover, $u^{\prime}(r)\geq 0$ and $v^{\prime}(r)\geq 0$ for all $r>0$.
Theorem 1.2 (Nonemptiness and closedness of $S$).
Let $S$ be the set of all central values $(a,b)\in[0,\infty)^{2}$ for which (1) admits an entire positive radial solution. Then $S$ is nonempty and closed.
Theorem 1.3 (Largeness on the edge).
If $(a,b)\in E=\partial S\cap(0,\infty)^{2}$ and $(u,v)$ is an entire positive radial solution of (1) with $u(0)=a$, $v(0)=b$, then
$\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=\infty.$
Paper structure. Section 2 contains some preliminaries, including the radial formulation of the system, local existence results, and the Keller–Osserman transform, which will be used in the proofs. Section 3 proves the existence of entire positive radial solutions for a nonempty set of central values. Section 4 shows that the set of all central values of entire solutions is nonempty and closed. Section 5 establishes that solutions corresponding to boundary points of this set are large, i.e., both components blow up at infinity. Section 6 presents concluding remarks and possible extensions. An Appendix contains an auxiliary proposition on the integrability of compositions of the nonlinearities.
2 Preliminaries: radial formulation, local existence and Keller–Osserman transform
We begin by recalling that radial solutions of (1) satisfy the ODE system
$\left\{\begin{array}[]{l}u^{\prime\prime}+\frac{n-1}{r}u^{\prime}=p(r)\,g(v),\qquad r>0,\\ v^{\prime\prime}+\frac{n-1}{r}v^{\prime}=q(r)\,f(u),\qquad r>0,\\ u(0)=a,\quad v(0)=b,\quad u^{\prime}(0)=v^{\prime}(0)=0.\end{array}\right$. (4)
By integrating twice and using the regularity at $r=0$, we obtain the equivalent integral formulation:
$u(r)$ $=a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g\big{(}v(s)\big{)}\,ds\,dt,$ (5)
$v(r)$ $=b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f\big{(}u(s)\big{)}\,ds\,dt.$ (6)
Lemma 2.1 (Local existence [10]).
For any $a,b>0$ there exists $\rho=\rho(a,b)>0$ and a positive solution $(u,v)\in C^{2}([0,\rho))^{2}$ of (4) with the given central values. Moreover, $u^{\prime}(r)\geq 0$ and $v^{\prime}(r)\geq 0$ for all $r\in(0,\rho)$.
Proof.
We work in the Banach space $C[0,\rho]\times C[0,\rho]$ with the norm
$\|(u,v)\|_{\infty}:=\max\{\|u\|_{\infty},\|v\|_{\infty}\}.$
Define the operator
$T:C[0,\rho]\times C[0,\rho]\rightarrow C[0,\rho]\times C[0,\rho]$
by
$T(u,v)(r):=\left(a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(v(s))\,ds\,dt,\ b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\,dt\right).$
Let
$X:=\left\{(u,v)\in C[0,\rho]\times C[0,\rho]\ \middle|\ \|(u,v)-(a,b)\|_{\infty}\leq\min\{a,b\}\right\}.$
For $\rho>0$ sufficiently small, the continuity of $p$, $q$, $f$, $g$ ensures that $T(X)\subset X$.  The mapping $T$ is continuous (by dominated convergence) and compact (by Arzela-Ascoli, since the image of $X$ is uniformly bounded and equicontinuous). By Schauder’s fixed point theorem, $T$ has a fixed point $(u,v)\in X$, which is a $C^{2}$ solution of (4) on $0,\rho)$.
Finally, $p,q\geq 0$ and $f,g\geq 0$ imply $u^{\prime},v^{\prime}$ are nondecreasing with $u^{\prime}(0)=v^{\prime}(0)=0$, hence $u^{\prime},v^{\prime}\geq 0$ on $(0,\rho)$. ∎
We define the maximal existence radius
[…]
Lemma 2.2 (Alternative [10]).
Let $a,b>0$. If $R_{a,b}<\infty$, then
$\lim_{r\uparrow R_{a,b}}u(r)=\lim_{r\uparrow R_{a,b}}v(r)=\infty.$
Proof.
Suppose, for contradiction, that $u$ remains bounded as $r\uparrow R_{a,b}$. Then $f(u(r))$ is bounded, say by $M>0$. From (4), we have
$v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)\leq q(r)M.$
Since $q$ is locally integrable, the right-hand side is locally bounded, so $v^{\prime}$ is locally Lipschitz and hence bounded on $0,R_{a,b})$. Integrating, $v$ is bounded on $0,R_{a,b})$. Then $g(v)$ is bounded, and the same argument applied to the $u$-equation shows $u^{\prime}$ is bounded. Thus $(u,v)$ and their derivatives are bounded up to $R_{a,b}$, so the local existence lemma allows extension beyond $R_{a,b}$, contradicting maximality. Therefore both $u$ and $v$ must blow up as $r\uparrow R_{a,b}$. ∎
Next, we note that (2) entails
$L_{f}:=\int_{1}^{\infty}\frac{ds}{g(f(s))\,}<\infty,\quad L_{g}:=\int_{1}^{\infty}\frac{ds}{f(g(s))\,}<\infty,$ (7)
for further details, see [2, 14]. We now introduce the Keller–Osserman-type transforms
$\Phi(t):=\int_{t}^{\infty}\frac{ds}{g(f(s))},\qquad\Psi(t):=\int_{t}^{\infty}\frac{ds}{f(g(s))}.$
By the finite reciprocal integral conditions (2), these integrals converge for all $t\geq 1$. Moreover, $\Phi$ and $\Psi$ extend to $C^{1}$ functions on $(0,\infty)$, are strictly decreasing, and satisfy
$\Phi^{\prime}(t)=-\frac{1}{g(f(t))}<0,\qquad\Psi^{\prime}(t)=-\frac{1}{f(g(t))}<0,$ (8)
together with the convexity properties
$\Phi^{\prime\prime}(t)=\frac{g^{\prime}(f(t))\,f^{\prime}(t)}{[g(f(t))]^{2}}>0,\qquad\Psi^{\prime\prime}(t)=\frac{f^{\prime}(g(t))\,g^{\prime}(t)}{[f(g(t))]^{2}}>0,$
where the inequalities follow from the assumptions $f^{\prime},g^{\prime}\geq 0$ and $f,g>0$ on $(0,\infty)$.
Lemma 2.3 (Forcing inequality).
Let $(u,v)$ be any positive radial $C^{2}$-solution of (4). Then, for every $r>0$, one has
$\left\{\begin{array}[]{c}u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)\ \leq\ p(r)\,g(f(u(r)))g(\frac{b}{f(a)}+\mathcal{L}_{q})\\ v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)\ \leq q(r)\,f(g(v(r)))f(\frac{a}{g(b)}+\mathcal{L}_{p}).\end{array}\right$. (9)
and consequently
$\Phi^{\prime\prime}(u)+\frac{n-1}{r}\Phi^{\prime}(u)\ \geq\ -p(r)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right),\qquad\Psi^{\prime\prime}(v)+\frac{n-1}{r}\Psi^{\prime}(v)\ \geq\ -q(r)\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right).$ (10)
Proof.
We start from the radial form of the system (4):
$u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)=p(r)\,g(v(r)),\qquad v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)=q(r)\,f(u(r)).$
From the integral representation (5)–(6) and the monotonicity of $u$ and $v$ (Lemma 2.1), we have for all $r>0$:
$v(r)\leq b+\mathcal{L}_{q}\,f(u(r)),$
where $\mathcal{L}_{q}$ is the finite constant defined in (3), i.e.
$\mathcal{L}_{q}:=\frac{1}{n-2}\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}q(\tau)\,d\tau\right)ds.$
Since $g$ is nondecreasing and positive, we can estimate
$g(v(r))\leq g\!\left(b+\mathcal{L}_{q}\,f(u(r))\right).$
By the multiplicative subadditivity of $g$,
$g\!\left(b+\mathcal{L}_{q}\,f(u(r))\right)\leq g\!\left(f(u(r))\right)\cdot g\!\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right).$
Moreover, since $u(r)\geq a$ for all $r\geq 0$ by monotonicity, we have
$\frac{b}{f(u(r))}+\mathcal{L}_{q}\ \leq\ \frac{b}{f(a)}+\mathcal{L}_{q},$
and thus
$g(v(r))\ \leq\ g(f(u(r)))\cdot g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
Substituting into the $u$-equation yields the first inequality in (9):
$u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)\ \leq\ p(r)\,g(f(u(r)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
A completely analogous argument, starting from
$u(r)\leq a+\mathcal{L}_{p}\,g(v(r)),$
and using the multiplicative subadditivity of $f$, gives
$v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)\ \leq\ q(r)\,f(g(v(r)))\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right),$
which is the second inequality in (9).
We now derive (10). By the chain rule,
$\Phi^{\prime}(u)=-\frac{u^{\prime}}{g(f(u))},$
and differentiating again,
$\Phi^{\prime\prime}(u)=-\frac{u^{\prime\prime}}{g(f(u))}+\frac{g^{\prime}(f(u))\,f^{\prime}(u)}{[g(f(u))]^{2}}\,(u^{\prime})^{2}.$
Since $f^{\prime},g^{\prime}\geq 0$, the second term is nonnegative, hence
$\Phi^{\prime\prime}(u)\ \geq\ -\frac{u^{\prime\prime}}{g(f(u))}.$
Adding $\frac{n-1}{r}\Phi^{\prime}(u)=-\frac{n-1}{r}\frac{u^{\prime}}{g(f(u))}$, we obtain
$\Phi^{\prime\prime}(u)+\frac{n-1}{r}\Phi^{\prime}(u)\ \geq\ -\frac{u^{\prime\prime}+\frac{n-1}{r}u^{\prime}}{g(f(u))}.$
Using the first inequality in (9) and cancelling $g(f(u))$ in numerator and denominator, we deduce
$\Phi^{\prime\prime}(u)+\frac{n-1}{r}\Phi^{\prime}(u)\ \geq\ -p(r)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
Similarly, for $\Psi$ we have
$\Psi^{\prime}(v)=-\frac{v^{\prime}}{f(g(v))},\qquad\Psi^{\prime\prime}(v)=-\frac{v^{\prime\prime}}{f(g(v))}+\frac{f^{\prime}(g(v))\,g^{\prime}(v)}{[f(g(v))]^{2}}\,(v^{\prime})^{2}\ \geq\ -\frac{v^{\prime\prime}}{f(g(v))}.$
Adding $\frac{n-1}{r}\Psi^{\prime}(v)$ and using the second inequality in (9) yields
$\Psi^{\prime\prime}(v)+\frac{n-1}{r}\Psi^{\prime}(v)\ \geq\ -q(r)\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right).$
This completes the proof. ∎
3 Existence of entire positive solutions
Let $c,d\in(0,\infty)$ be fixed. We define the set of admissible central values by
$\mathcal{T}:=\{(a,b)\in(0,\infty)^{2}\ :\ a<c\ \text{and}\ b<d\}.$
Consider the auxiliary scalar radial problem
$\left\{\begin{array}[]{l}z_{1}^{\prime\prime}+\frac{n-1}{r}z_{1}^{\prime}=p(r)g(\frac{b}{f(a)}+\mathcal{L}_{q})\,g\left(f\left(z_{1}\right)\right),\\ z_{2}^{\prime\prime}+\frac{n-1}{r}z_{2}^{\prime}=q(r)f(\frac{a}{g(b)}+\mathcal{L}_{p})f\left(g\left(z_{2}\right)\right),\\ z_{i}^{\prime}(0)=0,\ z_{1}(0)=c>0\text{, }z_{2}(0)=d>0.\end{array}\right$. (11)
where $\mathcal{L}_{p},\mathcal{L}_{q}$ are the finite constants
$\mathcal{L}_{p}:=\frac{1}{n-2}\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}p(\tau)\,d\tau\right)ds,\quad\mathcal{L}_{q}:=\frac{1}{n-2}\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}q(\tau)\,d\tau\right)ds,$
which are finite by the decay assumptions (3).
By the decay assumptions (3) and the Keller-Osserman-type condition inherited from (2), there exists a positive entire radial solution $\left(z_{1},z_{2}\right)$ of (11) which is large (i.e. $z_{1}(r)\rightarrow\infty$ and $z_{2}(r)\rightarrow\infty$, as $r\rightarrow\infty$) see [9] or [2, 14].
Proof of Theorem 1.1.
Let $(a,b)\in\mathcal{T}$  be fixed. By Lemma 2.1, there exists $\rho>0$ and a unique local positive radial solution $(u,v)\in C^{2}([0,\rho))^{2}$ of (1) with
$u(0)=a,\quad v(0)=b,\quad u^{\prime}(0)=v^{\prime}(0)=0.$
Define the maximal existence radius
$R:=\sup\left\{\rho>0\ :\ (u,v)\ \text{exists and is positive on }[0,\rho]\right\}.$
If $R=\infty$, then $(u,v)$ is already an entire positive radial solution and we are done. Suppose, for contradiction, that $R<\infty$.
Step 1: Comparison with $\left(z_{1},z_{2}\right)$. We claim that
$u(r)<z_{1}(r)\quad\text{and}\quad v(r)<z_{2}(r),\quad\forall r\in[0,R].$ (12)
Initial ordering. At $r=0$ we have
$u(0)=a<c=z_{1}(0),\quad v(0)=b<d=z_{2}(0).$
Thus the inequalities in (12) hold at $r=0$.
Definition of $R_{0}$. Let
$R_{0}:=\sup\left\{\rho\in(0,R]\ :\ u(r)<z_{1}(r)\ \text{and}\ v(r)<z_{2}(r)\ \forall r\in[0,\rho]\right\}.$
Clearly $R_{0}>0$ by continuity.
If $R_{0}=R$, then (12) is proved. Suppose instead that $R_{0}<R$.
Integral formulation and comparison. For $r\in[0,R_{0}]$, the integral form of (1) gives(5)-(6). Since $v(s)<z_{2}(s)$ for $s\leq R_{0}$ and $g$ is increasing, we have
$g(v(s))<g(z_{2}(s)).$
Moreover, by the structure of (11) and the constants $\mathcal{L}_{p},\mathcal{L}_{q}$, we can bound
$g(v(s))\leq g\!\left(f(u(s))\right)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right),$
and similarly for $f(u(s))$. Thus, at $r=R_{0}$,
$u(R_{0})$ $=a+\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}p(s)\,g(v(s))\,ds\right)dt$
$\leq a+g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}p(s)\,g(f(u(s)))\,ds\right)dt$
$<c+g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}p(s)\,g(f(z_{1}(s)))\,ds\right)dt$
$=z_{1}(R_{0}),$
where the strict inequality follows from $u(s)<z_{1}(s)$ for $s<R_{0}$ and the monotonicity of $g\circ f$.
Similarly,
$v(R_{0})$ $=b+\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\right)dt$
$\leq b+f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}q(s)\,f(g(z_{2}(s)))\,ds\right)dt$
$<d+f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}q(s)\,f(g(z_{2}(s)))\,ds\right)dt$
$=z_{2}(R_{0}).$
Thus
$u(R_{0})<z_{1}(R_{0}),\quad v(R_{0})<z_{2}(R_{0}).$
By continuity, there exists $\delta>0$ such that the inequalities persist on $0,R_{0}+\delta]$, contradicting the definition of $R_{0}$. Thus $R_{0}=R$ and (12) holds.
Step 2: Extension beyond $R$. From (12) and the fact that $z_{1}$, $z_{2}$ are finite on $[0,R]$, we have $u(R^{-})<\infty$ and $v(R^{-})<\infty$. By monotonicity of $u,v$ (Lemma 2.1), $u,v$ and their derivatives are bounded on $[0,R)$. Standard continuation for ODEs then extends $(u,v)$ beyond $R$, contradicting the definition of $R$. Hence $R=\infty$ and $(u,v)$ is an entire positive radial solution. ∎
4 The set of central values: Nonemptiness and closedness
We define the set of central values for which the system admits an entire positive radial solution as
[…]
Proof of Theorem 1.2.
We give the proof in:
Step 1: Nonemptiness. By Theorem 1.1, under our standing assumptions, there exists a nonempty set $T\subset(0,\infty)^{2}$ of admissible central values for which (1) has an entire positive radial solution. Since $T\subset S$, we conclude $S\neq\varnothing$.
Step 2: Closedness. Let $(a_{k},b_{k})\in S$ be a sequence converging to $(a_{0},b_{0})\in[0,\infty)^{2}$. For each $k$, let $(u_{k},v_{k})$ denote the corresponding entire positive radial solution of (5)-(6) defined by
$u_{k}(r)$ $=a_{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g\big{(}v_{k}(s)\big{)}\,ds\,dt,$
$v_{k}(r)$ $=b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f\big{(}u_{k}(s)\big{)}\,ds\,dt.$
with
$u_{k}(0)=a_{k},\quad v_{k}(0)=b_{k},\quad u_{k}^{\prime}(0)=v_{k}^{\prime}(0)=0.$
Uniform bounds on compact intervals. Fix $m\in\mathbb{N}$. Consider the scalar blow-up comparison problem (cf. (11) in Section 3):
$\begin{cases}z_{1}^{\prime\prime}(r)+\frac{n-1}{r}z_{1}^{\prime}(r)=p(r)g(\frac{b}{f(a)}+\mathcal{L}_{q})\,g\left(f\left(z_{1}\left(r\right)\right)\right),\\ z_{1}(0)=c>0,\quad z_{1}^{\prime}(0)=0,\end{cases}$
and similarly for $z_{2}$ with $q$ and $f(g(\cdot))$. By the Keller-Osserman-type conditions and the integrability of $p,q$, there exists $R_{m}>m$ such that $z_{1m}$ and $z_{2m}$ (the blow-up solutions with suitably large initial data) are finite on $[0,m]$ and blow up at $r=R_{m}$.
By the monotonicity of $u_{k}$ and $v_{k}$ in $r$, and the comparison principle for the radial ODEs, we have
$u_{k}(r)\leq z_{1m}(r),\quad v_{k}(r)\leq z_{2m}(r),\qquad\forall r\in[0,m],\ \forall k.$
Thus $\{u_{k}\}$ and $\{v_{k}\}$ are uniformly bounded on $[0,m]$.
$C^{1}$-bounds. From the radial form of (1) defined above:
$u_{k}^{\prime}\left(r\right)=r^{1-n}\int_{0}^{r}s^{n-1}p(s)\,g(v_{k}(s))\,ds,$
and the uniform bound on $v_{k}$ together with assumption on $p$ yield a uniform bound on $u_{k}^{\prime}$ on $[0,m]$. Similarly for $v_{k}^{\prime}$. Hence $\{u_{k}\}$ and $\{v_{k}\}$ are uniformly bounded in $C^{1}([0,m])$.
Compactness and passage to the limit. By the Arzel-Ascoli theorem, for each $m$ there exists a subsequence (still denoted by $u_{k},v_{k}$) converging in $C^{1}([0,m])$ to some $(u^{(m)},v^{(m)})$. A diagonal extraction yields a pair $(u,v)$ such that
$(u_{k},v_{k})\rightarrow(u,v)\quad\text{in }C_{\mathrm{loc}}^{1}([0,\infty)).$
Passing to the limit in the integral formulation of (1) (valid for radial $C^{2}$-solutions), we find that $(u,v)$ is an entire positive radial solution with
$u(0)=a_{0},\quad v(0)=b_{0}.$
Conclusion. Since $(u,v)$ is entire and positive, $(a_{0},b_{0})\in S$. Thus $S$ contains all its limit points and is therefore closed. ∎
5 Largeness on the edge
Proof of Theorem 1.3.
Let $(a,b)\in E$ and $(u,v)$ be an entire solution with these central values. From Theorem 1.2 we have $\partial S\subset S$. Choose a sequence $(a_{k},b_{k})\notin S$ with $(a_{k},b_{k})\rightarrow(a,b)$ and $a_{k},b_{k}>0$. By Lemma 2.1, for each $k$ there is a maximal radius $R_{k}<\infty$ for the solution $(u_{k},v_{k})$ with $u_{k}(0)=a_{k}$, $v_{k}(0)=b_{k}$. By Lemma 2.2,
$\lim_{r\rightarrow R_{k}^{-}}u_{k}(r)=\lim_{r\rightarrow R_{k}^{-}}v_{k}(r)=\infty.$
If $\{R_{k}\}$ were bounded, a subsequence would converge to some $R<\infty$, and the local uniform convergence $(u_{k},v_{k})\rightarrow(u,v)$ would imply $(u,v)$ blows up at $r=R$, contradicting that $(u,v)$ is entire. Hence $R_{k}\rightarrow\infty$.
Fix $r>0$; for $k$ large enough we have $r<R_{k}$ and, by monotonicity in the initial data, $u_{k}(r)\leq u(r)$ and $v_{k}(r)\leq v(r)$. By Lemma 2.3 we have
$\Phi^{\prime\prime}(u_{k}(r))+\frac{n-1}{r}\Phi^{\prime}(u_{k}(r))\ \geq\ -p(r)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right),$ (13)
and
$\Psi^{\prime\prime}(v_{k}(r))+\frac{n-1}{r}\Psi^{\prime}(v_{k}(r))\ \geq\ -q(r)\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right).$ (14)
Equations (13)–(14) can be rewritten as
$\left\{\begin{array}[]{c}\left(r^{n-1}\Phi^{\prime}(u_{k}\left(r\right))\right)^{\prime}\geq-r^{n-1}p(r)g(\frac{b}{f(a)}+\mathcal{L}_{q}),\\ \left(r^{n-1}\Psi^{\prime}(v_{k}\left(r\right))\right)^{\prime}\geq-r^{n-1}q(r)f(\frac{a}{g(b)}+\mathcal{L}_{p}).\end{array}\right$.
Integrating over $[0,r]$ with $0<r<R_{k}$ gives
$\left\{\begin{array}[]{c}\Phi^{\prime}(u_{k}\left(r\right))\geq-r^{1-n}\int_{0}^{r}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})ds,\\ \Psi^{\prime}(v_{k}\left(r\right))\geq-r^{1-n}\int_{0}^{r}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})ds.\end{array}\right$.
Next, integrate over $r,R_{k}]$. Since
$u_{k}(t)\rightarrow\infty\ \text{and}\ v_{k}(t)\rightarrow\infty\quad\text{as }t\rightarrow R_{k}^{-},$
we have
$\Phi(u_{k}(t))\rightarrow 0,\quad\Psi(v_{k}(t))\rightarrow 0\quad\text{as }t\rightarrow R_{k}^{-}.$
Thus
$\left\{\begin{array}[]{c}-\Phi(u_{k}\left(r\right))\geq-\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})dsdt\\ -\Psi(v_{k}\left(r\right))\geq-\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})dsdt\end{array}\right$.
or equivalently,
$\left\{\begin{array}[]{c}\Phi(u_{k}\left(r\right))\leq\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})dsdt,\\ \Psi(v_{k}\left(r\right))\leq\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})dsdt.\end{array}\right$.
Since $\Phi$ and $\Psi$ are strictly decreasing (see (8)), we can invert them to obtain
$\left\{\begin{array}[]{c}u_{k}\left(r\right)\geq\Phi^{-1}\left(\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})dsdt\right),\\ v_{k}\left(r\right)\geq\Psi^{-1}\left(\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})dsdt\right).\end{array}\right$.
Letting $k\rightarrow\infty$ gives $R_{k}\rightarrow R$ and $(u_{k},v_{k})\rightarrow(u,v)$ locally uniformly, hence
$\left\{\begin{array}[]{c}u\left(r\right)\geq\Phi^{-1}\left(\int_{r}^{R}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})dsdt\right),\\ v\left(r\right)\geq\Psi^{-1}\left(\int_{r}^{R}t^{1-n}\int_{0}^{t}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})dsdt\right).\end{array}\right$.
Finally, letting $r\rightarrow R$ and using
$\lim_{t\rightarrow 0^{+}}\Phi^{-1}(t)=\lim_{t\rightarrow 0^{+}}\Psi^{-1}(t)=\infty,$
we deduce
$\lim_{r\rightarrow R}u(r)=\lim_{r\rightarrow R}v(r)=\infty.$
Since $(u,v)$ is entire with $(a,b)\in E\subset S$, we must have $R=\infty$. Thus $u(r),v(r)\rightarrow\infty$ as $r\rightarrow\infty$, i.e., $(u,v)$ is an entire large solution. ∎
Remark 2 (Criticality).
The Keller–Osserman condition (2) is sharp for this method: if it fails, the barrier function is infinite at infinity and the radial integration argument cannot force blow-up (see our work [3] for this case).
Remark 3.
Theorem 1.3 shows that the “edge” $\partial S$ of the set of admissible central values corresponds precisely to the threshold between bounded entire solutions and large entire solutions. This is analogous to the scalar case treated in the classical Keller–Osserman theory, but here the coupling between $u$ and $v$ requires a more delicate comparison argument.
6 Concluding remarks
We have established:
  • Existence of entire positive radial solutions for all central values $(a,b)$ in a nonempty open set $\mathcal{T}$.
  • The set $S$ of all such central values is closed in $0,\infty)^{2}$.
  • Points on the edge $E$ correspond to large solutions, where both components blow up at infinity.
The key novelty lies in the use of the Keller–Osserman-type transforms $\Phi$ and $\Psi$ together with the forcing inequalities of Lemma 2.3, which allow us to control the growth of solutions under the finite reciprocal integral conditions (2). This framework unifies and extends previous results for scalar equations and special systems, and applies to a broad class of nonlinearities and weight functions.