### v2

# Existence and large radial solutions for an elliptic system under finite new Keller-Osserman integral conditions

Abstract
We investigate the existence, multiplicity, and asymptotic behavior of entire positive radial solutions to the semilinear elliptic system
$\Delta u=p(|x|)\,g(v),\qquad\Delta v=q(|x|)\,f(u),\qquad x\in\mathbb{R}^{n},\ n\geq 3,$
under new Keller–Osserman-type integral conditions on the nonlinearities $f$ and $g$, and decay constraints on the radial weight functions $p$ and $q$. The nonlinearities are assumed continuous on $0,\infty)$, differentiable on $(0,\infty)$, nondecreasing, multiplicatively subadditive, vanishing at the origin, and strictly positive elsewhere, satisfying the finite reciprocal integral conditions
[…]
The radial weights satisfy $\int_{0}^{\infty}s\,p(s)\,ds<\infty$ and $\int_{0}^{\infty}s\,q(s)\,ds<\infty$, with $\min(p,q)$ not compactly supported. Within this framework, we establish three main results: (i) existence of infinitely many entire positive radial solutions for each central value $(a,b)$ in a nonempty open set $\mathcal{T}\subset(0,\infty)^{2}$; (ii) closedness of the set $S$ of all admissible central values in $0,\infty)^{2}$; and (iii) largeness (blow-up at infinity) of solutions corresponding to boundary points $E=\partial S\cap(0,\infty)^{2}$. The proofs employ novel comparison techniques with auxiliary scalar problems, Arzelã -Ascoli compactness arguments, and a subharmonic functional approach tailored to the reciprocal integral conditions via Keller–Osserman-type transforms. Our results extend classical Keller–Osserman theory to a broad class of coupled elliptic systems with general nonlinearities and weight functions, unifying and generalizing several existing frameworks.
MSC 2020: 35J47 (Primary), 35B40, 35B09, 35J60, 35B33 (Secondary).
Keywords: Semilinear elliptic systems; Keller–Osserman condition; entire positive radial solutions; large solutions; blow-up at infinity; multiplicative subadditivity; comparison principle; weighted elliptic systems; Arzelã -Ascoli theorem.
1 Introduction
The study of positive entire solutions to nonlinear elliptic systems has a long and rich history, driven both by deep theoretical challenges in nonlinear analysis and by diverse applications in physics, geometry, and biology. A classical starting point is the scalar equation
$\Delta u=f(u),$
for which Keller [5] and Osserman [10] established the celebrated Keller–Osserman condition, providing sharp integral criteria for the existence of large solutions (see also [12] for more details). These foundational results have inspired extensive research on systems of elliptic equations, where the interaction between components introduces new analytical difficulties.
For semilinear elliptic systems, Lair and Wood [8] proved the existence of entire large positive solutions under suitable growth and integrability conditions, while Peterson-Wood [11] and Li-Yang [9] extended the analysis to non-monotone systems, but with superlinear nonlinearities. More recently, Lair [7] investigated competitive-type systems, obtaining sharp existence results for large solutions.
In addition to the aforementioned works, the contributions of [1, 3, 4, 13] have played a significant role in advancing the theory of elliptic systems, particularly in elucidating the crucial interplay between nonlinearity and weight functions.
In this work, we consider the semilinear elliptic system
$\Delta u=p(r)\,g(v),\qquad\Delta v=q(r)\,f(u),\qquad r=|x|,\quad x\in\mathbb{R}^{n},\ n\geq 3,$ (1)
under the following assumptions:
  • (F1)
    $f,g:[0,\infty)\rightarrow[0,\infty)$ are continuous on $0,\infty)$, differentiable on $(0,\infty)$, non-decreasing, with $f(0)=g(0)=0$ and $f(s),g(s)>0$ for all $s>0$;
  • (F2)
    multiplicative subadditivity:
    $f(sr)\leq f(s)f(r),\quad g(lt)\leq g(l)g(t),\quad\forall s,r,l,t\geq 0;$
  • (F3)
    finite Keller–Osserman-type integrals:
    $L_{f}:=\int_{1}^{\infty}\frac{dt}{\sqrt{\int_{0}^{t}g(f(z))\,dz}}<\infty,\quad L_{g}:=\int_{1}^{\infty}\frac{dt}{\sqrt{\int_{0}^{t}f(g(z))\,dz}}<\infty;$ (2)
  • (PQ)
    $p,q:[0,\infty)\rightarrow[0,\infty)$ are continuous, not both identically zero, and satisfy
    $\mathcal{L}_{p}$ $=\frac{1}{n-2}\int_{0}^{\infty}s\,p(s)\,ds=\lim_{r\rightarrow\infty}\mathcal{P}(r)<\infty,$
    $\mathcal{L}_{q}$ $=\frac{1}{n-2}\int_{0}^{\infty}s\,q(s)\,ds=\lim_{r\rightarrow\infty}\mathcal{Q}(r)<\infty,$ (3)
    where
    $\mathcal{P}(r)=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt,\quad\mathcal{Q}(r)=\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt,$
    and $\min(p,q)$ does not have compact support.
We seek positive radial $C^{2}$-solutions $(u(r),v(r))$ of (1) with prescribed central values
$u(0)=a>0,\quad v(0)=b>0,\quad u^{\prime}(0)=v^{\prime}(0)=0.$
A positive entire large solution is a pair $(u,v)$ of $C^{2}$ radial functions such that $u(r),v(r)>0$ for all $r\geq 0$ and
$\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=\infty.$ (4)
In [7], Lair proved existence of such solutions for power-type nonlinearities $f(u)=u^{\theta}$, $g(v)=v^{\sigma}$ under certain integrability conditions on $p$ and $q$. Our goal is to extend this to general nonlinearities satisfying (2).
Remark 1.1 (Relation to Lair’s result).
When $g(v)=v^{\sigma}$ and $f(u)=u^{\theta}$ with $\sigma,\theta>0$ and $\sigma\cdot\theta>1$, the Keller–Osserman integrals in (2) converge, recovering the power-growth case in [7].
Our main objectives are:
  • to prove the existence of infinitely many positive entire radial solutions for a nonempty set of central values;
  • to show that the set of admissible central values is closed;
  • to establish that solutions corresponding to boundary points of this set are large, i.e., both components blow up at infinity.
The novelty of our approach lies in the construction of a new subharmonic functional adapted to the finite reciprocal integral conditions on $f$ and $g$, which allows us to complete, extend and unify the frameworks developed in [3, 4, 1, 5, 10, 8, 7, 11] to a broader class of nonlinearities and weight functions.
We now state our three main results.
Theorem 1.2 (Existence).
For every $(a,b)$ in a suitable nonempty set of admissible central values $\mathcal{T}\subset(0,\infty)^{2}$, there exists an entire positive radial solution $(u,v)\in C^{2}([0,\infty))^{2}$ of (1) with $u(0)=a$, $v(0)=b$. Moreover, $u^{\prime}(r)\geq 0$ and $v^{\prime}(r)\geq 0$ for all $r>0$.
Theorem 1.3 (Nonemptiness and closedness of $S$).
Let $S$ be the set of all central values $(a,b)\in(0,\infty)^{2}$ for which (1) admits an entire positive radial solution. Then $S$ is nonempty and closed.
Theorem 1.4 (Largeness on the edge).
If $(a,b)\in E=\partial S\cap(0,\infty)^{2}$ and $(u,v)$ is an entire positive radial solution of (1) with $u(0)=a$, $v(0)=b$, then
$\lim_{r\rightarrow\infty}u(r)=\lim_{r\rightarrow\infty}v(r)=\infty.$
Paper structure. Section 2 contains some preliminaries, including the radial formulation of the system, local existence results, and the Keller–Osserman transform, which will be used in the proofs. Section 3 proves the existence of entire positive radial solutions for a nonempty set of central values. Section 4 shows that the set of all central values of entire solutions is nonempty and closed. Section 5 establishes that solutions corresponding to boundary points of this set are large, i.e., both components blow up at infinity. Compatibility between Theorem 1.4 and Theorem 2.4 of [3] is given in Section 6. Section 7 presents concluding remarks and possible extensions. An Appendix contains an auxiliary proposition on the integrability of compositions of the nonlinearities.
2 Preliminaries: radial formulation, local existence and Keller–Osserman transform
We begin by recalling that radial solutions of (1) satisfy the ODE system
$\left\{\begin{array}[]{l}u^{\prime\prime}+\frac{n-1}{r}u^{\prime}=p(r)\,g(v),\qquad r>0,\\ v^{\prime\prime}+\frac{n-1}{r}v^{\prime}=q(r)\,f(u),\qquad r>0,\\ u(0)=a,\quad v(0)=b,\quad u^{\prime}(0)=v^{\prime}(0)=0.\end{array}\right$. (5)
By integrating twice and using the regularity at $r=0$, we obtain the equivalent integral formulation:
$u(r)$ $=a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g\big(v(s)\big)\,ds\,dt,$ (6)
$v(r)$ $=b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f\big(u(s)\big)\,ds\,dt.$ (7)
Before proving our results, we note first that for $a,b\in\left(0,\infty\right)$ any solution of the system of integral equations (6)–(7) valid for all $r\geq 0$ will also be a positive entire solution of (1) but not necessarily (4).
Lemma 2.1 (Local existence [11]).
For any $a,b>0$ there exists $\rho=\rho(a,b)>0$ and a positive solution $(u,v)\in C^{2}([0,\rho))^{2}$ of (5) with the given central values. Moreover, $u^{\prime}(r)\geq 0$ and $v^{\prime}(r)\geq 0$ for all $r\in[0,\rho)$.
Proof.
We work in the Banach space $C[0,\rho]\times C[0,\rho]$ with the norm
$\|(u,v)\|_{\infty}:=\max\{\|u\|_{\infty},\|v\|_{\infty}\}.$
Define the operator
$T:C[0,\rho]\times C[0,\rho]\rightarrow C[0,\rho]\times C[0,\rho]$
by
$T(u,v)(r):=\left(a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(v(s))\,ds\,dt,\ b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\,dt\right).$
Let
$X:=\left\{(u,v)\in C[0,\rho]\times C[0,\rho]\ \middle|\ \|(u,v)-(a,b)\|_{\infty}\leq\min\{a,b\}\right\}.$
For $\rho>0$ sufficiently small, the continuity of $p$, $q$, $f$, $g$ ensures that $T(X)\subset X$. The mapping $T$ is continuous (by dominated convergence) and compact (by Arzela-Ascoli, since the image of $X$ is uniformly bounded and equicontinuous). By Schauder’s fixed point theorem, $T$ has a fixed point $(u,v)\in X$, which is a $C^{2}$ solution of (5) on $0,\rho)$.
Finally, $p,q\geq 0$ and $f,g\geq 0$ imply $u^{\prime},v^{\prime}$ are nondecreasing with $u^{\prime}(0)=v^{\prime}(0)=0$, hence $u^{\prime},v^{\prime}\geq 0$ on $0,\rho)$. ∎
We define the maximal existence radius
[…]
Lemma 2.2 (Alternative [11]).
Let $a,b>0$. If $R_{a,b}<\infty$, then
$\lim_{r\uparrow R_{a,b}}u(r)=\lim_{r\uparrow R_{a,b}}v(r)=\infty.$
Proof.
Suppose, for contradiction, that $u$ remains bounded as $r\uparrow R_{a,b}$. Then $f(u(r))$ is bounded, say by $M>0$. From (5), we have
$v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)\leq q(r)M.$
Since $q$ is locally integrable, the right-hand side is locally bounded, so $v^{\prime}$ is locally Lipschitz and hence bounded on $0,R_{a,b})$. Integrating, $v$ is bounded on $0,R_{a,b})$. Then $g(v)$ is bounded, and the same argument applied to the $u$-equation shows $u^{\prime}$ is bounded. Thus $(u,v)$ and their derivatives are bounded up to $R_{a,b}$, so the local existence lemma allows extension beyond $R_{a,b}$, contradicting maximality. Therefore both $u$ and $v$ must blow up as $r\uparrow R_{a,b}$. ∎
Next, we note that (2) entails
$L_{f}:=\int_{1}^{\infty}\frac{ds}{g(f(s))\,}<\infty,\quad L_{g}:=\int_{1}^{\infty}\frac{ds}{f(g(s))\,}<\infty,$ (8)
for further details, see [2, 14]. We now introduce the Keller–Osserman-type transforms
$\Phi(t):=\int_{t}^{\infty}\frac{ds}{g(f(s))},\qquad\Psi(t):=\int_{t}^{\infty}\frac{ds}{f(g(s))}.$
By the finite reciprocal integral conditions (2), these integrals converge for all $t\geq 1$. Moreover, $\Phi$ and $\Psi$ extend to $C^{1}$ functions on $(0,\infty)$, are strictly decreasing, and satisfy
$\Phi^{\prime}(t)=-\frac{1}{g(f(t))}<0,\qquad\Psi^{\prime}(t)=-\frac{1}{f(g(t))}<0,$ (9)
together with the convexity properties
$\Phi^{\prime\prime}(t)=\frac{g^{\prime}(f(t))\,f^{\prime}(t)}{[g(f(t))]^{2}}>0,\qquad\Psi^{\prime\prime}(t)=\frac{f^{\prime}(g(t))\,g^{\prime}(t)}{[f(g(t))]^{2}}>0,$
where the inequalities follow from the assumptions $f^{\prime},g^{\prime}\geq 0$ and $f,g>0$ on $(0,\infty)$.
Lemma 2.3 (Forcing inequality).
Let $a,b>0$ and $(u,v)$ be any positive radial $C^{2}$-solution of (5). Then, for every $r>0$, one has
$\left\{\begin{array}[]{c}u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)\ \leq\ p(r)\,g(f(u(r)))g(\frac{b}{f(a)}+\mathcal{L}_{q})\\ v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)\ \leq q(r)\,f(g(v(r)))f(\frac{a}{g(b)}+\mathcal{L}_{p}).\end{array}\right$. (10)
and consequently
$\Phi^{\prime\prime}(u)+\frac{n-1}{r}\Phi^{\prime}(u)\ \geq\ -p(r)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right),\qquad\Psi^{\prime\prime}(v)+\frac{n-1}{r}\Psi^{\prime}(v)\ \geq\ -q(r)\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right).$ (11)
Proof.
We start from the radial form of the system (5):
$u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)=p(r)\,g(v(r)),\qquad v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)=q(r)\,f(u(r)).$
From the integral representation (6)–(7) and the monotonicity of $u$ and $v$ (Lemma 2.1), we have for all $r>0$:
$v(r)=b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\,dt.$
Since $u(s)\geq u(0)=a>0$ for all $s\geq 0$ by monotonicity, and $f$ is nondecreasing, we obtain
$v(r)\leq b+f(u(r))\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt.$
By the decay assumption (3), the double integral converges:
$\mathcal{L}_{q}:=\frac{1}{n-2}\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}q(\tau)\,d\tau\right)ds=\lim_{r\to\infty}\mathcal{Q}(r)<\infty.$
Therefore, for all $r\geq 0$,
$v(r)\leq b+\mathcal{L}_{q}\,f(u(r)).$ (12)
Since $g$ is nondecreasing and positive, we estimate
$g(v(r))\leq g\!\left(b+\mathcal{L}_{q}\,f(u(r))\right).$
Now we apply the multiplicative subadditivity property (F2). Writing
$b+\mathcal{L}_{q}\,f(u(r))=f(u(r))\cdot\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right),$
and noting that $f(u(r))>0$ for $u(r)>0$, we invoke (F2) with $l=f(u(r))$ and $t=\frac{b}{f(u(r))}+\mathcal{L}_{q}$ to obtain
$g\!\left(b+\mathcal{L}_{q}\,f(u(r))\right)=g\!\left(f(u(r))\cdot\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right)\right)\leq g(f(u(r)))\cdot g\!\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right).$
Since $u(r)\geq a$ for all $r\geq 0$ by monotonicity, and $f$ is nondecreasing, we have $f(u(r))\geq f(a)$, hence
$\frac{b}{f(u(r))}\leq\frac{b}{f(a)},$
and thus by monotonicity of $g$,
$g\!\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right)\leq g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
Combining these estimates, we obtain
$g(v(r))\ \leq\ g(f(u(r)))\cdot g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
Substituting into the $u$-equation yields the first inequality in (10):
$u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)=p(r)g(v(r))\ \leq\ p(r)\,g(f(u(r)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
By complete symmetry, we establish the bound for $u$:
$u(r)\leq a+\mathcal{L}_{p}\,g(v(r)),$
where $\mathcal{L}_{p}$ is defined in (3). Applying the multiplicative subadditivity of $f$ and monotonicity of $f$, $v(r)\geq b$ implies
$f(u(r))\leq f(a+\mathcal{L}_{p}g(v(r)))\leq f(g(v(r)))f\left(\frac{a}{g(v(r))}+\mathcal{L}_{p}\right)\leq f(g(v(r)))f\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right),$
yielding the second inequality in (10):
$v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)\ \leq\ q(r)\,f(g(v(r)))\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right).$
We now derive (11) from (10). Recall that $\Phi(t)=\int_{t}^{\infty}\frac{ds}{g(f(s))}$ is $C^{1}$ on $(0,\infty)$ with
$\Phi^{\prime}(t)=-\frac{1}{g(f(t))}<0.$
Applying the chain rule to $\Phi\circ u$, we obtain
$\frac{d}{dr}\Phi(u(r))=\Phi^{\prime}(u(r))\,u^{\prime}(r)=-\frac{u^{\prime}(r)}{g(f(u(r)))},$
and differentiating a second time,
$\frac{d^{2}}{dr^{2}}\Phi(u(r))$ $=\Phi^{\prime\prime}(u(r))(u^{\prime}(r))^{2}+\Phi^{\prime}(u(r))u^{\prime\prime}(r)$
$=-\frac{u^{\prime\prime}(r)}{g(f(u(r)))}+\frac{d}{dt}\left[-\frac{1}{g(f(t))}\right]_{t=u(r)}(u^{\prime}(r))^{2}$
$=-\frac{u^{\prime\prime}(r)}{g(f(u(r)))}+\frac{g^{\prime}(f(u(r)))\,f^{\prime}(u(r))}{[g(f(u(r)))]^{2}}\,(u^{\prime}(r))^{2}.$
By assumption (F1), $f^{\prime},g^{\prime}\geq 0$ and $u^{\prime}\geq 0$, so the second term is nonnegative. Hence
$\Phi^{\prime\prime}(u(r))\ \geq\ -\frac{u^{\prime\prime}(r)}{g(f(u(r)))}.$
Moreover,
$\frac{n-1}{r}\Phi^{\prime}(u(r))=-\frac{n-1}{r}\,\frac{u^{\prime}(r)}{g(f(u(r)))}.$
Adding these two inequalities gives
$\Phi^{\prime\prime}(u(r))+\frac{n-1}{r}\Phi^{\prime}(u(r))\ \geq\ -\frac{u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)}{g(f(u(r)))}.$ (13)
From the first inequality in (10), we have
$u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)\leq p(r)\,g(f(u(r)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
Substituting this into (13) and noting that $g(f(u(r)))$ cancels in the numerator and denominator, we deduce
$\Phi^{\prime\prime}(u(r))+\frac{n-1}{r}\Phi^{\prime}(u(r))\ \geq\ -\frac{p(r)\,g(f(u(r)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)}{g(f(u(r)))}=-p(r)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
This is the first inequality in (11).
By complete symmetry, we have for $\Psi(t)=\int_{t}^{\infty}\frac{ds}{f(g(s))}$:
$\Psi^{\prime}(v(r))=-\frac{v^{\prime}(r)}{f(g(v(r)))},$
and
$\Psi^{\prime\prime}(v(r))=-\frac{v^{\prime\prime}(r)}{f(g(v(r)))}+\frac{f^{\prime}(g(v(r)))\,g^{\prime}(v(r))}{[f(g(v(r)))]^{2}}\,(v^{\prime}(r))^{2}\ \geq\ -\frac{v^{\prime\prime}(r)}{f(g(v(r)))}.$
Thus
$\Psi^{\prime\prime}(v(r))+\frac{n-1}{r}\Psi^{\prime}(v(r))\ \geq\ -\frac{v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)}{f(g(v(r)))}.$
Using the second inequality in (10) and canceling $f(g(v(r)))$ yields
$\Psi^{\prime\prime}(v(r))+\frac{n-1}{r}\Psi^{\prime}(v(r))\ \geq\ -q(r)\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right),$
which is the second inequality in (11). This completes the proof. ∎
3 Existence of entire positive solutions
Let $c,d\in(0,\infty)$ be fixed. We define the set of admissible central values by
$\mathcal{T}:=\{(a,b)\in(0,\infty)^{2}\ :\ a<c\ \text{and}\ b<d\}.$
Consider the auxiliary scalar radial problem
$\left\{\begin{array}[]{l}z_{1}^{\prime\prime}+\frac{n-1}{r}z_{1}^{\prime}=p(r)g(\frac{b}{f(a)}+\mathcal{L}_{q})\,g\left(f\left(z_{1}\right)\right),\\ z_{2}^{\prime\prime}+\frac{n-1}{r}z_{2}^{\prime}=q(r)f(\frac{a}{g(b)}+\mathcal{L}_{p})f\left(g\left(z_{2}\right)\right),\\ z_{i}^{\prime}(0)=0,\ z_{1}(0)=c>0\text{, }z_{2}(0)=d>0.\end{array}\right$. (14)
where $\mathcal{L}_{p},\mathcal{L}_{q}$ are the finite constants
$\mathcal{L}_{p}:=\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}p(\tau)\,d\tau\right)ds,\quad\mathcal{L}_{q}:=\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}q(\tau)\,d\tau\right)ds,$
which are finite by the decay assumptions (3).
By the decay assumptions (3) and the Keller-Osserman-type condition inherited from (2), there exists a positive entire radial solution $\left(z_{1},z_{2}\right)$ of (14) which is large (i.e. $z_{1}(r)\rightarrow\infty$ and $z_{2}(r)\rightarrow\infty$, as $r\rightarrow\infty$) see [6] or [2, 14].
Proof of Theorem 1.2.
Let $(a,b)\in\mathcal{T}$ be fixed, i.e., $0<a<c$ and $0<b<d$. By Lemma 2.1, there exists $\rho>0$ and a unique local positive radial solution $(u,v)\in C^{2}([0,\rho))^{2}$ of (1) satisfying the integral equations (6)–(7) with initial values
$u(0)=a,\quad v(0)=b,\quad u^{\prime}(0)=v^{\prime}(0)=0.$
Moreover, by Lemma 2.1, $u^{\prime}(r)\geq 0$ and $v^{\prime}(r)\geq 0$ for all $r\in[0,\rho)$, so $u$ and $v$ are nondecreasing.
Define the maximal existence radius
$R:=\sup\left\{\rho>0\ :\ (u,v)\ \text{exists as a positive $C^{2}$ solution on }[0,\rho]\right\}.$
By Lemma 2.1, $R>0$. If $R=\infty$, then $(u,v)$ is already an entire positive radial solution and we are done.
We now prove that $R=\infty$ by contradiction. Suppose that $R<\infty$.
Step 1: Comparison with $\left(z_{1},z_{2}\right)$. We claim that
$u(r)<z_{1}(r)\quad\text{and}\quad v(r)<z_{2}(r),\quad\forall r\in[0,R].$ (15)
Initial ordering. At $r=0$ we have
$u(0)=a<c=z_{1}(0),\quad v(0)=b<d=z_{2}(0).$
Thus the inequalities in (15) hold at $r=0$.
Definition of $R_{0}$. Let
$R_{0}:=\sup\left\{\rho\in(0,R]\ :\ u(r)<z_{1}(r)\ \text{and}\ v(r)<z_{2}(r)\text{,}\ \forall r\in[0,\rho]\right\}.$
Clearly $R_{0}>0$ by continuity.
If $R_{0}=R$, then (15) is proved. Suppose instead that $R_{0}<R$.
Integral formulation and comparison. For $r\in[0,R_{0}]$, we have from the integral formulation (6)–(7):
$u(r)$ $=a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(v(s))\,ds\,dt,$
$v(r)$ $=b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\,dt.$
By Lemma 2.3, specifically inequality (10), we have for all $s\in[0,R_{0}]$:
$g(v(s))\leq g(f(u(s)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
Therefore, for any $r\in[0,R_{0}]$,
$u(r)$ $=a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(v(s))\,ds\,dt$
$\leq a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(f(u(s)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\,ds\,dt.$ (16)
Recall that $z_{1}$ satisfies (see (14)):
$z_{1}(r)=c+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(\frac{b}{f(a)}+\mathcal{L}_{q})\,g(f(z_{1}(s)))\,ds\,dt.$
Since $u(s)<z_{1}(s)$ for all $s\in[0,R_{0})$ and $g\circ f$ is nondecreasing (by (F1)), we have
$g(f(u(s)))<g(f(z_{1}(s)))\quad\text{for all }s\in(0,R_{0}).$
Moreover, $a<c$ by assumption. Therefore, for $r\in(0,R_{0}]$, from (16):
$u(r)$ $\leq a+g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(f(u(s)))\,ds\,dt$
$<c+g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(f(z_{1}(s)))\,ds\,dt$
$=z_{1}(r),$
where the strict inequality uses $a<c$ and $g(f(u(s)))<g(f(z_{1}(s)))$ for $s\in(0,r]$. In particular, taking $r=R_{0}$, we obtain
$u(R_{0})<z_{1}(R_{0}).$
By exactly the same reasoning, using Lemma 2.3 to obtain
$f(u(s))\leq f(g(v(s)))\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right),$
and noting that $v(s)<z_{2}(s)$ for $s\in[0,R_{0})$ implies $f(g(v(s)))<f(g(z_{2}(s)))$, and $b<d$, we deduce
$v(r)$ $=b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\,dt$
$\leq b+f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(g(v(s)))\,ds\,dt$
$<d+f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(g(z_{2}(s)))\,ds\,dt$
$=z_{2}(r).$
In particular, at $r=R_{0}$,
$v(R_{0})<z_{2}(R_{0}).$
Thus we have shown that
$u(R_{0})<z_{1}(R_{0})\quad\text{and}\quad v(R_{0})<z_{2}(R_{0}).$
By the continuity of $u,v,z_{1},z_{2}$, there exists $\delta>0$ such that
$u(r)<z_{1}(r)\quad\text{and}\quad v(r)<z_{2}(r)\quad\text{for all }r\in[0,R_{0}+\delta],$
contradicting the supremum definition of $R_{0}$. Therefore, we must have $R_{0}=R$ and (15) holds for all $r\in[0,R]$.
Step 2: Extension beyond $R$. From (15), we have
$u(r)\leq z_{1}(r)\quad\text{and}\quad v(r)\leq z_{2}(r)\quad\text{for all }r\in[0,R).$
Since $z_{1}$ and $z_{2}$ are entire positive radial solutions (hence finite on $[0,R]$ for any finite $R$), there exist constants $M_{1},M_{2}>0$ such that
$z_{1}(r)\leq M_{1}\quad\text{and}\quad z_{2}(r)\leq M_{2}\quad\text{for all }r\in[0,R].$
Therefore, by comparison,
$u(r)\leq M_{1}\quad\text{and}\quad v(r)\leq M_{2}\quad\text{for all }r\in[0,R).$
Since $u$ and $v$ are nondecreasing (Lemma 2.1), the limits
$\lim_{r\to R^{-}}u(r)=:u(R^{-})\leq M_{1}<\infty,\quad\lim_{r\to R^{-}}v(r)=:v(R^{-})\leq M_{2}<\infty$
exist and are finite.
Moreover, from the radial equations (5),
$u^{\prime}(r)=r^{1-n}\int_{0}^{r}s^{n-1}p(s)g(v(s))\,ds,$
and since $v(s)\leq M_{2}$ for $s\in[0,R)$ and $g$ is continuous, we have $g(v(s))\leq g(M_{2})$ for all $s\in[0,R)$. Thus
$u^{\prime}(r)\leq r^{1-n}\int_{0}^{r}s^{n-1}p(s)g(M_{2})\,ds\leq g(M_{2})\mathcal{P}(R)<\infty,$
where $\mathcal{P}(R)<\infty$ by assumption (PQ). Similarly, $v^{\prime}(r)\leq f(M_{1})\mathcal{Q}(R)<\infty$ for all $r\in[0,R)$.
Therefore, $(u,v)$ and $(u^{\prime},v^{\prime})$ are uniformly bounded on $[0,R)$. By the fundamental theorem of calculus and uniform continuity, the limits
$\lim_{r\to R^{-}}(u(r),u^{\prime}(r),v(r),v^{\prime}(r))$
exist and are finite. By the standard existence and uniqueness theory for ODEs (Lemma 2.1 applied with initial data at $r=R$), the solution can be extended beyond $R$ to some interval $0,R+\varepsilon)$ with $\varepsilon>0$. This contradicts the maximality of $R$.
Hence, we conclude that $R=\infty$, and $(u,v)$ is an entire positive radial solution of (1) with $u(0)=a$, $v(0)=b$, and $u^{\prime}(r)\geq 0$, $v^{\prime}(r)\geq 0$ for all $r\geq 0$. ∎
4 The set of central values: Nonemptiness and closedness
We define the set of central values for which the system admits an entire positive radial solution as
[…]
Proof of Theorem 1.3.
We give the proof in:
Step 1: Nonemptiness. By Theorem 1.2, under our standing assumptions, there exists a nonempty set $T\subset(0,\infty)^{2}$ of admissible central values for which (1) has an entire positive radial solution. Since $T\subset S$, we conclude $S\neq\varnothing$.
Step 2: Closedness. Let $(a_{k},b_{k})\in S$ be a sequence converging to $(a_{0},b_{0})\in[0,\infty)^{2}$. For each $k$, let $(u_{k},v_{k})$ denote the corresponding entire positive radial solution of (6)-(7) defined by
$u_{k}(r)$ $=a_{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g\big(v_{k}(s)\big)\,ds\,dt,$
$v_{k}(r)$ $=b_{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f\big(u_{k}(s)\big)\,ds\,dt.$
with
$u_{k}(0)=a_{k},\quad v_{k}(0)=b_{k},\quad u_{k}^{\prime}(0)=v_{k}^{\prime}(0)=0.$
Uniform bounds on compact intervals. Fix $m\in\mathbb{N}$ arbitrarily. Since $(a_{k},b_{k})\to(a_{0},b_{0})$, the sequence $\{(a_{k},b_{k})\}$ is bounded. Without loss of generality, assume $a_{k}\leq A$ and $b_{k}\leq B$ for all $k$ and some constants $A,B>0$.
For each $k$, the solution $(u_{k},v_{k})$ satisfies, by the proof of Theorem 1.2, a comparison with auxiliary solutions $(z_{1,k},z_{2,k})$ solving:
$\begin{cases}z_{1,k}^{\prime\prime}+\frac{n-1}{r}z_{1,k}^{\prime}=p(r)g(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q})\,g\left(f\left(z_{1,k}\right)\right),\\ z_{1,k}(0)=c_{k}>a_{k},\quad z_{1,k}^{\prime}(0)=0,\end{cases}$ (17)
and similarly for $z_{2,k}$. However, to obtain uniform bounds independent of $k$, we construct a single comparison solution that dominates all $(u_{k},v_{k})$ simultaneously.
Let $\tilde{a}=\sup_{k}a_{k}\leq A$ and $\tilde{b}=\sup_{k}b_{k}\leq B$ (both finite by boundedness). Choose constants $\bar{c}>\tilde{a}$ and $\bar{d}>\tilde{b}$. Consider the auxiliary scalar problems:
$\begin{cases}\bar{z}_{1}^{\prime\prime}+\frac{n-1}{r}\bar{z}_{1}^{\prime}=p(r)g(\frac{B}{f(\inf_{k}a_{k})}+\mathcal{L}_{q})\,g\left(f\left(\bar{z}_{1}\right)\right),\\ \bar{z}_{1}(0)=\bar{c},\quad\bar{z}_{1}^{\prime}(0)=0,\end{cases}$
and
$\begin{cases}\bar{z}_{2}^{\prime\prime}+\frac{n-1}{r}\bar{z}_{2}^{\prime}=q(r)f(\frac{A}{g(\inf_{k}b_{k})}+\mathcal{L}_{p})f\left(g\left(\bar{z}_{2}\right)\right),\\ \bar{z}_{2}(0)=\bar{d},\quad\bar{z}_{2}^{\prime}(0)=0.\end{cases}$
By the Keller-Osserman-type conditions (F3) and the decay assumptions (PQ), these problems admit entire positive radial solutions $\bar{z}_{1}$ and $\bar{z}_{2}$ (see [6, 2]).
By the comparison principle (as in the proof of Theorem 1.2), since $a_{k}<\bar{c}$ and $b_{k}<\bar{d}$ for all $k$, and the right-hand sides of the equations for $(u_{k},v_{k})$ are dominated by those for $(\bar{z}_{1},\bar{z}_{2})$ (due to monotonicity of $f,g$ and the choices of constants), we obtain
$u_{k}(r)\leq\bar{z}_{1}(r),\quad v_{k}(r)\leq\bar{z}_{2}(r),\qquad\forall r\geq 0,\ \forall k\in\mathbb{N}.$
In particular, for the fixed $m$, we have uniform bounds:
$u_{k}(r)\leq\bar{z}_{1}(m)=:M_{1}(m)<\infty,\quad v_{k}(r)\leq\bar{z}_{2}(m)=:M_{2}(m)<\infty,\quad\forall r\in[0,m],\ \forall k.$
Thus $\{u_{k}\}$ and $\{v_{k}\}$ are uniformly bounded on $[0,m]$.
$C^{1}$-bounds. From the integral representation (6), we have
$u_{k}^{\prime}\left(r\right)=r^{1-n}\int_{0}^{r}s^{n-1}p(s)\,g(v_{k}(s))\,ds,\quad r\in(0,m].$
Since $v_{k}(s)\leq M_{2}(m)$ for all $s\in[0,m]$ and $g$ is continuous and nondecreasing, we have
$g(v_{k}(s))\leq g(M_{2}(m))=:G_{m}<\infty,\quad\forall s\in[0,m],\ \forall k.$
Thus
$u_{k}^{\prime}(r)\leq r^{1-n}\int_{0}^{r}s^{n-1}p(s)G_{m}\,ds\leq G_{m}\mathcal{P}(m)<\infty,\quad\forall r\in[0,m],\ \forall k,$
where $\mathcal{P}(m)<\infty$ by assumption (PQ). By symmetry,
$v_{k}^{\prime}(r)\leq f(M_{1}(m))\mathcal{Q}(m)<\infty,\quad\forall r\in[0,m],\ \forall k.$
Hence $\{u_{k}\}$ and $\{v_{k}\}$ are uniformly bounded in $C^{1}([0,m])$.
Moreover, from the equations (5), for $r\in[\delta,m]$ with $\delta>0$,
$u_{k}^{\prime\prime}(r)=-\frac{n-1}{r}u_{k}^{\prime}(r)+p(r)g(v_{k}(r))\leq\frac{n-1}{\delta}G_{m}\mathcal{P}(m)+\|p\|_{L^{\infty}([0,m])}G_{m},$
so $\{u_{k}^{\prime\prime}\}$ is uniformly bounded on $\delta,m]$. Similarly for $v_{k}^{\prime\prime}$. Thus $\{(u_{k},v_{k})\}$ is equicontinuous in $C^{1}$ on $\delta,m]$ for any $\delta>0$.
Compactness and passage to the limit. By the Arzelà-Ascoli theorem, for each $m\in\mathbb{N}$ there exists a subsequence (denote it by $\{u_{k_{j}^{(m)}},v_{k_{j}^{(m)}}\}$) converging in $C^{1}([0,m])$ to some $(u^{(m)},v^{(m)})\in C^{1}([0,m])$. By a standard diagonal argument, we extract a subsequence $\{k_{\ell}\}_{\ell=1}^{\infty}$ (still denoted by $\{k\}$ for simplicity) such that
$(u_{k},v_{k})\rightarrow(u,v)\quad\text{in }C_{\mathrm{loc}}^{1}([0,\infty)),$
i.e., $(u_{k},v_{k})\to(u,v)$ in $C^{1}([0,m])$ for every $m\in\mathbb{N}$.
We now show that $(u,v)$ is a solution of (1). Passing to the limit in the integral equations (6)–(7), using the continuity of $f,g,p,q$ and the dominated convergence theorem (justified by the uniform bounds), we obtain for every $r\geq 0$:
$u(r)$ $=\lim_{k\to\infty}u_{k}(r)=\lim_{k\to\infty}\left[a_{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(v_{k}(s))\,ds\,dt\right]$
$=a_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(v(s))\,ds\,dt,$
and similarly for $v$. Thus $(u,v)$ satisfies (6)–(7), hence is a $C^{2}$ positive radial solution of (1) with
$u(0)=a_{0},\quad v(0)=b_{0},\quad u^{\prime}(0)=v^{\prime}(0)=0.$
Conclusion. Since $(u,v)$ is entire and positive, $(a_{0},b_{0})\in S$. Thus $S$ contains all its limit points and is therefore closed. ∎
5 Largeness on the edge
Proof of Theorem 1.4.
Let $(a,b)\in E=\partial S\cap(0,\infty)^{2}$ be a boundary point of $S$ in the positive quadrant, and let $(u,v)$ be an entire positive radial solution of (1) with $u(0)=a$, $v(0)=b$.
Since $S$ is closed (Theorem 1.3), we have $\partial S\subset S$, so $(a,b)\in S$ and the solution $(u,v)$ exists and is entire.
Step 1: Approximating sequence. Since $(a,b)\in\partial S$, there exists a sequence $(a_{k},b_{k})\notin S$ with $(a_{k},b_{k})\rightarrow(a,b)$ as $k\to\infty$, and $a_{k},b_{k}>0$ for all $k$.
For each $k$, by Lemma 2.1, there exists a maximal radius $R_{k}\in(0,\infty]$ and a positive radial solution $(u_{k},v_{k})\in C^{2}([0,R_{k}))^{2}$ of (1) with $u_{k}(0)=a_{k}$, $v_{k}(0)=b_{k}$, and $u_{k}^{\prime}(0)=v_{k}^{\prime}(0)=0$.
Since $(a_{k},b_{k})\notin S$, the solution $(u_{k},v_{k})$ cannot be extended to all of $0,\infty)$, hence $R_{k}<\infty$ for all $k$. By Lemma 2.2,
$\lim_{r\rightarrow R_{k}^{-}}u_{k}(r)=\lim_{r\rightarrow R_{k}^{-}}v_{k}(r)=\infty.$ (18)
Step 2: The sequence $\{R_{k}\}$ is unbounded. Suppose, for contradiction, that $\{R_{k}\}$ is bounded. Then there exists $R_{\star}<\infty$ such that $R_{k}\leq R_{\star}$ for all $k$. Passing to a subsequence if necessary, we may assume $R_{k}\to R_{0}\leq R_{\star}$ as $k\to\infty$.
[…]
But $(u,v)$ is the unique solution with these initial values (by uniqueness in Lemma 2.1), so $\tilde{u}=u$ and $\tilde{v}=v$ on $0,R_{0})$. Thus $u_{k}\to u$ and $v_{k}\to v$ locally uniformly on $0,R_{0})$.
However, from (18), $u_{k}(r)\to\infty$ as $r\to R_{k}^{-}\to R_{0}^{-}$, which would imply $u(r)\to\infty$ as $r\to R_{0}^{-}$, contradicting that $(u,v)$ is entire. Therefore, $\{R_{k}\}$ is unbounded, i.e., $R_{k}\rightarrow\infty$ as $k\to\infty$.
Step 3: Application of the Keller-Osserman transforms. Now that we have established $R_{k}\to\infty$, we apply the transforms $\Phi$ and $\Psi$ to derive lower bounds on $u$ and $v$.
Fix $r>0$ arbitrarily. For $k$ sufficiently large, we have $r<R_{k}$, so $(u_{k},v_{k})$ is well-defined on $[0,r]$.
By Lemma 2.3, specifically inequality (11), the functions $u_{k}$ and $v_{k}$ satisfy for all $\rho\in(0,R_{k})$:
$\Phi^{\prime\prime}(u_{k}(\rho))+\frac{n-1}{\rho}\Phi^{\prime}(u_{k}(\rho))\ \geq\ -p(\rho)\,g\!\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right),$ (19)
and
$\Psi^{\prime\prime}(v_{k}(\rho))+\frac{n-1}{\rho}\Psi^{\prime}(v_{k}(\rho))\ \geq\ -q(\rho)\,f\!\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right).$ (20)
Since $(a_{k},b_{k})\to(a,b)$ and $f,g$ are continuous, the constants $\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\to\frac{b}{f(a)}+\mathcal{L}_{q}$ and $\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\to\frac{a}{g(b)}+\mathcal{L}_{p}$ as $k\to\infty$. For $k$ sufficiently large, these constants are bounded, say by $C_{1}$ and $C_{2}$ respectively.
Inequalities (19)–(20) can be rewritten in divergence form:
$\left\{\begin{array}[]{c}\left(\rho^{n-1}\Phi^{\prime}(u_{k}(\rho))\right)^{\prime}\geq-\rho^{n-1}p(\rho)g(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}),\\ \left(\rho^{n-1}\Psi^{\prime}(v_{k}(\rho))\right)^{\prime}\geq-\rho^{n-1}q(\rho)f(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}).\end{array}\right$. (21)
Step 4: Integration and derivation of lower bounds. Integrating (21) over $[0,r]$ with $0<r<R_{k}$, and using $\Phi^{\prime}(u_{k}(0))=-\frac{u_{k}^{\prime}(0)}{g(f(u_{k}(0)))}=0$ (since $u_{k}^{\prime}(0)=0$), we obtain
$r^{n-1}\Phi^{\prime}(u_{k}(r))\geq-\int_{0}^{r}s^{n-1}p(s)g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)ds.$ (22)
Dividing by $r^{n-1}$ yields
$\Phi^{\prime}(u_{k}(r))\geq-r^{1-n}\int_{0}^{r}s^{n-1}p(s)g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)ds.$
Similarly,
$\Psi^{\prime}(v_{k}(r))\geq-r^{1-n}\int_{0}^{r}s^{n-1}q(s)f\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right)ds.$
Now, we integrate over $r,R_{k}]$. Since $u_{k}(t)\rightarrow\infty$ and $v_{k}(t)\rightarrow\infty$ as $t\rightarrow R_{k}^{-}$ (by (18)), and since $\Phi(t)=\int_{t}^{\infty}\frac{ds}{g(f(s))}\to 0$ as $t\to\infty$ (by the finite integral condition (8)), we have
$\lim_{t\to R_{k}^{-}}\Phi(u_{k}(t))=0,\quad\lim_{t\to R_{k}^{-}}\Psi(v_{k}(t))=0.$ (23)
Integrating the inequality for $\Phi^{\prime}(u_{k}(\rho))$ over $\rho\in[r,R_{k})$, we obtain
$0-\Phi(u_{k}(r))$ $=\lim_{t\to R_{k}^{-}}\Phi(u_{k}(t))-\Phi(u_{k}(r))=\int_{r}^{R_{k}}\Phi^{\prime}(u_{k}(\rho))\,d\rho$
$\geq-\int_{r}^{R_{k}}\rho^{1-n}\int_{0}^{\rho}s^{n-1}p(s)g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)ds\,d\rho.$
Thus
$\Phi(u_{k}(r))\leq g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt.$ (24)
Similarly,
$\Psi(v_{k}(r))\leq f\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right)\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt.$ (25)
Step 5: Inversion and passage to the limit. Since $\Phi$ and $\Psi$ are strictly decreasing (see (9)), they are invertible. From (24)–(25), we obtain
$u_{k}(r)$ $\geq\Phi^{-1}\left(g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt\right),$ (26)
$v_{k}(r)$ $\geq\Psi^{-1}\left(f\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right)\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt\right).$ (27)
Now, let $k\to\infty$. We have:
  • $R_{k}\to\infty$ (Step 2);
  • $(a_{k},b_{k})\to(a,b)$, so $g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)\to g\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)=:G_{0}$ and $f\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right)\to f\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)=:F_{0}$ by continuity;
  • For any fixed $r>0$,
    $\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt\to\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt=\mathcal{P}(\infty)-\mathcal{P}(r)$
    as $k\to\infty$, where $\mathcal{P}(\infty)=\mathcal{L}_{p}(n-2)<\infty$ by assumption (PQ). Similarly for $q$.
By the continuity of $\Phi^{-1}$ and passing to the limit in (26) as $k\to\infty$, we obtain (noting that $u_{k}\to u$ locally uniformly from Step 2’s argument):
$u(r)\geq\Phi^{-1}\left(G_{0}\left[\mathcal{L}_{p}(n-2)-\mathcal{P}(r)\right]\right)=\Phi^{-1}\left(G_{0}\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt\right).$ (28)
Similarly,
$v(r)\geq\Psi^{-1}\left(F_{0}\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt\right).$ (29)
Step 6: Blow-up at infinity. Since $p,q$ are not both identically zero and $\min(p,q)$ does not have compact support (assumption (PQ)), at least one of the integrals $\int_{0}^{\infty}s^{n-1}p(s)\,ds$ or $\int_{0}^{\infty}s^{n-1}q(s)\,ds$ is positive. Thus, as $r\to\infty$,
$\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt\to 0\quad\text{and/or}\quad\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt\to 0.$
Since $\Phi(t)\to 0$ as $t\to\infty$ (by (8)), we have $\Phi^{-1}(s)\to\infty$ as $s\to 0^{+}$. Similarly for $\Psi^{-1}$. Therefore, from (28)–(29), as $r\to\infty$,
$u(r)\geq\Phi^{-1}(G_{0}\cdot[\text{positive term}\to 0])\to\infty,\quad v(r)\geq\Psi^{-1}(F_{0}\cdot[\text{positive term}\to 0])\to\infty.$
Hence,
$\lim_{r\to\infty}u(r)=\infty\quad\text{and}\quad\lim_{r\to\infty}v(r)=\infty,$
proving that $(u,v)$ is a large solution. This completes the proof. ∎
Remark 5.1 (Criticality).
The Keller–Osserman condition (2) is sharp for this method: if it fails, the barrier function is infinite at infinity and the radial integration argument cannot force blow-up (see our work [3] for this case).
Remark 5.2.
Theorem 1.4 shows that the “edge” $\partial S$ of the set of admissible central values corresponds precisely to the threshold between bounded entire solutions and large entire solutions. This is analogous to the scalar case treated in the classical Keller–Osserman theory, but here the coupling between $u$ and $v$ requires a more delicate comparison argument.
6 Compatibility between Theorem 1.4 and Theorem 2.4 of [3]
6.1 General framework
Both ELRS (Theorem 1.4) and Theorem 2.4 of [3] study radial elliptic systems of the form (1), where in [3] $S_{k_{i}}$ denotes the $k_{i}$-Hessian operator (the case $k_{i}=1$ being the Laplacian). Here $p,q$ are nonnegative radial weights and $f,g$ are positive nonlinearities.
  • In ELRS, finite Keller–Osserman-type conditions are imposed to ensure the existence of both bounded entire and large (blow-up at infinity) radial solutions, depending on the central values $(u(0),v(0))$.
  • In T2.4, the focus is on positive bounded entire radial solutions for the Hessian system with gradient term, under radial symmetry, a factorization structure for $f,g$, and a strict hierarchy between the functionals $\mathcal{P},\mathcal{Q},\Phi,\Psi$.
6.2 A common particular example
Let
$g(s)$ $=$ $s^{\alpha},\quad f(s)=s^{\beta},\quad\alpha=\beta=\frac{6}{5},\quad s\geq 0,$
$p(r)$ $=$ $q(r)=\frac{1}{(1+r)^{\gamma}},\quad\gamma=5,\quad r\geq 0.$
Then:
  • $f$ and $g$ are continuous, positive, strictly increasing on $(0,\infty)$;
  • the classical Keller–Osserman condition holds;
  • the potential integrals are finite since $p,q\sim r^{-5}$;
  • $\alpha\beta>1$ ensures the finiteness of the $\circ$-functionals;
  • the strict hierarchy $P_{i,j}(\infty)<H_{i,j}(\infty)$ required in T2.4 is satisfied.
6.3 Applicability of the two results
For ELRS:
The finite Keller–Osserman conditions together with finite potential integrals place this example in the class where there exist both bounded entire solutions (central values in the interior of $S$) and large solutions (central values on the boundary $E=\partial S$). ELRS distinguishes these regimes by the location of the central value.
For T2.4:
All hypotheses (radial symmetry, factorization, monotonicity, strict hierarchy) are met. The theorem yields a positive bounded entire radial solution for central values satisfying the strict inequality in the hypotheses. Such values lie in the interior of $S$ in the ELRS framework.
6.4 Why there is no contradiction
Let $S$ be the set of central values for which entire solutions exist, and $E=\partial S\cap(0,\infty)^{2}$ its “edge” in the positive quadrant.
  • T2.4 produces bounded entire solutions for $(u(0),v(0))$ in the interior of $S$.
  • ELRS asserts that for $(u(0),v(0))\in E$ any entire solution is large, while for interior points bounded entire solutions also exist.
In the above example, choosing central values according to T2.4 (strict inequality) places the solution in $\mathrm{int}(S)$, outside the scope of the “large solution” conclusion of ELRS. The two results are therefore compatible: they describe different behaviours (bounded vs. large) in disjoint regions of the central value space, even for the same coefficients.
Moreover, $E$ is not empty under the stated hypotheses: taking the boundary $\partial S$ in $\mathbb{R}^{2}$ and intersecting with $(0,\infty)^{2}$ yields a nonempty edge where the qualitative behaviour changes. Points of $E$ mark the transition between bounded entire solutions (interior) and large solutions (edge), making $E$ both topologically and analytically significant.
7 Concluding remarks
We have established:
  • Existence of entire positive radial solutions for all central values $(a,b)$ in a nonempty open set $\mathcal{T}$.
  • The set $S$ of all such central values is closed in $0,\infty)^{2}$.
  • Points on the edge $E$ correspond to large solutions, where both components blow up at infinity.
The key novelty lies in the use of the Keller–Osserman-type transforms $\Phi$ and $\Psi$ together with the forcing inequalities of Lemma 2.3, which allow us to control the growth of solutions under the finite reciprocal integral conditions (2). This framework unifies and extends previous results for scalar equations and special systems, and applies to a broad class of nonlinearities and weight functions.